# Free fermions, neutrality and modular transformations

### v1

@@ 1 Introduction
where $k$ (integer or half-integer) is called the weight, $a,b,c,d$ are integers, $\epsilon$ a complex number that depends on the $\mathrm{SL}(2,\mathbb{Z})$ element, and $\tau$ is a complex number typically taken to live on the upper-half plane $\mathbb{H}^{+}$. In recent years the class of modular forms has been extended significantly to functions, such as quasi-modular and quantum modular forms, for which the difference between the function and its modular transformation is not zero but exhibits “anomalous” behaviour, which can still have well-defined, desirable features. This has led to intense study of such more exotic modular objects and a growing appreciation of their physical relevance (see e.g. the lectures [1] for an introduction from a physics perspective, and references to the mathematical literature).
@@ 2 Motivation: Higher spin/$\mathcal{W}$-algebra correspondence
A very fruitful setting for studying the AdS/CFT correspondence is the duality between higher-spin gravity on $AdS_{3}$ and 2d minimal models with $\mathcal{W}$ algebra symmetry [12]. The duality is based on the asymptotic symmetry group of Vasiliev’s higher spin theory [13] being $\mathcal{W}_{\infty}[\lambda]$ [14, 15], with $\lambda$ a tunable parameter. One of the most successful tests of this duality is the matching of the spectrum [16]. In particular, [17] studied the thermodynamics of black holes with a spin-3 chemical potential and compared it to the result obtained from the CFT side. Although the duality should apply for more general CFT’s, there are special points in parameter space where the CFT computation is very straightforward, and here we will focus on the case of a free fermion CFT, where the Vasiliev parameter $\lambda=0$. For this case, it is straightforward to compute the partition function and obtain [17]:
@@
Now, one of the main results of [28] was that the $\zeta^{0}$ component of $\Theta(\zeta,q,w)$ can be expanded in powers of $w^{2n}$, the coefficients of which are quasi-modular forms of degree $6n$. This result played an important role in the counting of maps from genus-$g$ curves to genus 1 [30], where in that case $n=g-1$. In that case the underlying free-fermionic description [31] is half-integer moded, so the specific $\Theta$ function appearing is “$\theta_{3}$-like” and in particular its expression as a sum is over integer powers (see appendix A). The required $\Theta$ in our case (2.4) is “$\theta_{2}$-like” with half-integer powers in the sum, so (in either of the forms above) extracting the $\zeta^{0}$ terms is slightly more complicated. We will get a flavour of this difference in the next section, for the $\alpha=0$ case, and we will see that it will lead to an object with quantum- instead of quasi- modular properties.
Having explained our motivation, we will not discuss the higher-spin setting further in this work, and will instead focus on the much more modest goal of imposing neutrality in the $\alpha=0$ case, and only for one fermion, i.e. $D=1$. (The $D=2$ case is treated in appendix B).
@@ 3 Neutrality, Dyson’s crank, and partial theta functions
To see the high-temperature behaviour of $Z_{\text{charged}}$ we need to study it as $\tau\rightarrow i0^{+}$, where $q\rightarrow 1$ and the $q$-expansion is not very helpful. As is standard, we apply modular transformations (see appendix A) to write (3.1) as
@@
$Z_{\text{neutral}}(\tau)=\oint\frac{\mathrm{d}\zeta}{2\pi i\zeta}Z(u,\tau)=\frac{q^{-\frac{1}{12}}}{\eta(\tau)}\sum_{k>0,~{}k\text{odd}}(-1)^{\frac{k-1}{2}}q^{k^{2}/8}=\frac{q^{\frac{1}{24}}}{\eta(\tau)}\sum_{n\geq 0}(-1)^{n}q^{\frac{n(n+1)}{2}}\;,$ (3.9)
where we set $k=2n+1$ in the last step. This series has recently been shown to be a generating functions of Dyson’s non-negative crank partition function, see [34, 35], where it is given in the equivalent form
@@ 5 Modularity and high-temperature expansion
$\int_{0}^{\infty}f(x)~{}\eta(ix)\mathrm{d}x=\int_{0}^{\infty}\frac{F(t)}{\cosh\sqrt{\pi t}}\mathrm{d}t\;,$ (5.14)
@@
Together with the fact that $H(1,2;-1/\tau)$ is exponentially small as $\tau\rightarrow 0$ (due to the $\hat{q}^{1/8}$ factor, where $\hat{q}=e^{-2\pi i/\tau}$), we arrive at the small-$\tau$ expansion of $H(1,2;\tau$):
@@ 6 The canonical partition function
$\frac{1}{\pi^{2}}\frac{\partial^{2}\ln Z(u,\tau)}{\partial u^{2}}|_{u=0}=(2i\hat{\tau}/\pi+1/\cos(\pi u))|_{u=0}=2i\hat{\tau}/\pi+1$ (6.5)
@@
$Z_{C}=Z_{\text{charged}}(\tau)\sqrt{-i\tau}(1+\frac{\pi(-i\tau)}{4}+\cdots)\;,$ (6.7)
@@ 7 The asymptotic density of states
In this section we would like to find the asymptotic density of states of our neutral partition function, which is the main ingredient entering the Cardy formula. It is instructive to first review the computation for the full (charged) partition function, in order to highlight the differences with the neutral case. This computation is standard and reviews can be found in many sources, e.g. [5, 6, 53]. As discussed, when $\tau\rightarrow 0$, $q\rightarrow 1$ and it is convenient to re-express $Z_{\text{charged}}=Z(u,\tau)$ in terms of $\hat{\tau}=-1/\tau$, such that $\hat{q}\rightarrow 0$ (see (3.4)). We define a generating function $\rho(n)$ as
@@
$\rho_{\text{neutral}}^{(0)}=\frac{\pi}{2\sqrt{2}6^{3/4}(n+1/12)^{3/4}}\left(-i^{\frac{1}{2}}J_{3/2}(z)\right)\;.$ (7.8)
Expanding the Bessel function for large $n$ (see the Appendix), we find:
@@ 8 Discussion
Recalling our motivation from section 2, one is of course interested in how our results might apply to the higher-spin/$\mathcal{W}$ algebra correspondence. A first aspect to note is that for a comparison with the bulk one needs $D$ fermions, and although in that case the grand canonical partition function is just $Z(u,\tau)^{D}$, the restriction to the $\zeta^{0}$ part as in (3.8) is not immediate. In Appendix B we work it out for $D=2$, but ideally we would require an expression for large $D$ to make contact with the bulk. The generalisation, at $D=1$, of the non-negative crank partition function to include the spin-3 chemical potential, and which has the same relation to the partition function (2.1) as $Z_{\text{neutral}}$ does to (3.1), appears straightforward. However, one would then need to establish a suitable generalisation of the modular transformation (5.10) in order to make progress.
@@ Acknowledgements
We are thankful to Sam van Leuven for very enlightening discussions. MM was supported by a DST/NRF Innovation Postdoctoral Fellowship during the initial stages of this work, and by the National Institute for Theoretical and Computational Sciences (NITheCS) during the later stages.
@@ Appendix A Definitions
$\theta_{1}^{\prime}(u|\tau)=\frac{\mathrm{d}}{\mathrm{d}u}\theta_{1}(u|\tau)=2\pi q^{\frac{1}{8}}\sum_{k=0}(-1)^{k}(2k+1)\cos(\pi(2k+1)u)q^{\frac{k(k+1)}{2}}\;.$ (A.5)
@@
$\theta^{\prime}$ is also modular, of weight $3/2$. We will only need $\theta^{\prime}(0,\tau)$, whose modular transformation follows from the equality
$\theta^{\prime}(0,\tau)=2\pi\left(\eta(\tau)\right)^{3}\;.$ (A.9)
@@ Appendix B Free fermions
For completeness, in this appendix we record the relevant generators of $\mathcal{W}_{1+\infty}$ [19], in its fermionic realisation (see e.g. [55] for a review of the various realisations of $\mathcal{W}_{1+\infty}$ and [56] for more details on the computation of the character formula). More details can be found in [17], from whose presentation we differ just by simple rescalings to simplify the presentation. We consider a complex fermion with OPE
@@ Appendix C The $D=2$ case
As discussed after (3.8), the restriction to the neutral sector is expected to be more challenging for general $D$. To get a flavour of the complications that arise, in this appendix we derive the corresponding result for $D=2$.
@@
$Z_{\text{neutral}}^{D=2}=\oint\frac{\mathrm{d}\zeta}{2\pi i\zeta}(Z(u,\tau))^{2}=1+4q^{2}+8q^{3}+13q^{4}+24^{5}+46q^{6}+80q^{7}+139q^{9}+\cdots$ (C.1)
@@
$\begin{split}\theta_{2}^{2}(u,\tau)&=\frac{1}{2}\left(\theta_{3}(0,\tau/2)\theta_{3}(u,\tau/2)-\theta_{4}(0,\tau/2)\theta_{4}(u,\tau/2)\right)\\ &=\frac{1}{2}\left(\sum_{k\in\mathbb{Z}}q^{\frac{k^{2}}{4}}\sum_{\ell\in\mathbb{Z}}\zeta^{\ell}q^{\frac{\ell^{2}}{4}}-\sum_{k\in\mathbb{Z}}(-1)^{k}q^{\frac{k^{2}}{4}}\sum_{\ell\in\mathbb{Z}}(-1)^{\ell}\zeta^{\ell}q^{\frac{\ell^{2}}{4}}\right)\\ &=\frac{1}{2}\sum_{k\in\mathbb{Z}}\sum_{\ell\in\mathbb{Z}}\left(1-(-1)^{k+\ell}\right)\zeta^{\ell}q^{\frac{k^{2}+\ell^{2}}{4}}\\ &=\sum_{k,\ell\in\mathbb{Z},k+\ell\;\text{odd}}\zeta^{\ell}q^{\frac{k^{2}+\ell^{2}}{2}}\;.\end{split}$ (C.2)
@@
$\theta_{2}^{2}(u,\tau)=\sum_{n>0,n\;\text{odd}}(2+\zeta^{n}+\zeta^{-n})q^{\frac{n^{2}}{4}}+2\sum_{\begin{subarray}{c}m>0,n>m\\ m+n\;\text{odd}\end{subarray}}(\zeta^{m}+\zeta^{-m}+\zeta^{n}+\zeta^{-n})q^{\frac{m^{2}+n^{2}}{4}}\;.$ (C.3)
@@
$\begin{split}Z_{\text{neutral}}^{D=2}&=\frac{q^{-\frac{1}{4}}}{(q;q)_{\infty}^{2}}\oint\frac{\mathrm{d}\zeta}{2\pi i\zeta}\left(\frac{\theta_{2}(u,\tau)}{2\cos\pi u}\right)^{2}=\frac{1}{(q;q)_{\infty}^{2}}\sum_{\begin{subarray}{c}m\geq 0,n>m\\ m+n\;\text{odd}\end{subarray}}(2-\delta_{m,0})(-1)^{m}(n-m)q^{\frac{m^{2}+n^{2}-1}{4}}\\ &=\frac{1}{(q;q)_{\infty}^{2}}\left(1\!-\!2q\!+\!3q^{2}\!+\!2q^{3}\!-\!6q^{4}\!+\!3q^{6}\!+\!6q^{7}\!-\!10q^{9}\!+\!2q^{10}\!-\!6q^{11}\!+\!7q^{12}\!+\!10q^{13}+\cdots\right)\;.\end{split}$ (C.5)
@@
This slightly peculiar $q$-series agrees, after expanding the prefactor, with the empirical result (C.1). Although in this case we are not aware of any mathematical results regarding its modular properties, the physical setting is precisely the same as the $D=1$ case, so we can expect that (a) its leading behaviour for $\tau\rightarrow i0^{+}$ will be the same as the unconstrained partition function, i.e. $\sim e^{\frac{\pi i}{6\tau}-\frac{\pi i\tau}{3}}$ (just the square of the $D=1$ case), and more importantly (b) that $\hat{H}(\tau):=(q;q)_{\infty}^{2}Z_{\text{neutral}}^{D=2}$ has a well-defined expansion perturbatively in $\tau$, analogous to (5.3), corresponding to a quantum modular form of weight 1. Of course, the first statement more or less follows from the $1/(q;q)_{\infty}^{2}$ factors (expressed in terms of $\eta(\tau)^{2}$ and modular transformed), as the remainder is unlikely to alter the exponential behaviour. Evaluating the series for large $n$, we indeed find very good numerical agreement with the leading behaviour
$\rho_{\text{neutral}}^{D=2}=\frac{e^{\sqrt{\frac{2}{3}}\pi\sqrt{2(n+1/6)}}}{16\sqrt{3}(n+1/6)}\;.$ (C.6)
@@
$Z_{\text{neutral}}^{D=2}=\frac{e^{\frac{\pi i}{6\tau}-\frac{\pi i\tau}{3}}}{4\sqrt{2}}\sqrt{-i\tau}\left(1+\frac{\pi(-i\tau)}{4}\right)\;,$ (C.7)

### v2

@@ 1 Introduction
where $k$ (integer or half-integer) is called the weight, $a,b,c,d$ are integers, $\epsilon$ a complex number that depends on the $\mathrm{SL}(2,\mathbb{Z})$ element, and $\tau$ is a complex number taken to live on the upper-half plane $\mathbb{H}^{+}$. In recent years the class of modular forms has been extended significantly to functions, such as quasi-modular and quantum modular forms, for which the difference between the function and its modular transformation is not zero but exhibits “anomalous” behaviour, which can still have well-defined, desirable features. This has led to intense study of such more exotic modular objects and a growing appreciation of their physical relevance (see e.g. the lectures [1] for an introduction from a physics perspective, and references to the mathematical literature).
@@ 2 Motivation: Higher spin/$\mathcal{W}$-algebra correspondence
A very fruitful setting for studying the AdS/CFT correspondence is the duality between higher-spin gravity on $AdS_{3}$ and 2d minimal models with $\mathcal{W}$ algebra symmetry [12]. The duality is based on the asymptotic symmetry group of Vasiliev’s higher spin theory [13] being $\mathcal{W}_{\infty}[\lambda]$ [14, 15], with $\lambda$ a tunable parameter. One of the most successful tests of this duality is the matching of the spectrum [16]. In particular, [17] studied the thermodynamics of black holes with a spin-3 chemical potential and compared it to the result obtained from the CFT side. Although the duality should apply for more general CFT’s, there are special points in parameter space where the CFT computation is very straightforward, and here we will focus on the case of a free fermion CFT, where the Vasiliev parameter $\lambda=0$. For this case, it is straightforward to compute the partition function [17] to obtain, in the conventions of appendix B:
@@
Now, one of the main results of [28] was that the $\zeta^{0}$ component of $\Theta(\zeta,q,w)$ can be expanded in powers of $w^{2n}$, the coefficients of which are quasi-modular forms of degree $6n$. This result played an important role in the counting of maps from genus-$g$ curves to genus 1 [30], where in that case $n=g-1$. In that case the underlying free-fermionic description [31] is half-integer moded, so the specific $\Theta$ function appearing is “$\theta_{3}$-like” and in particular its expression is a sum is over integer powers (see appendix A). The required $\Theta$ in our case (2.4) is “$\theta_{2}$-like” with half-integer powers in the sum, so (in either of the forms above) extracting the $\zeta^{0}$ terms is slightly more delicate. We will get a flavour of this difference in the next section, for the $\alpha=0$ case, and we will see that it will lead to an object with quantum- instead of quasi- modular properties.
Having explained our motivation, we will not discuss the higher-spin setting further in this work, and will instead focus on the much more modest goal of imposing neutrality in the $\alpha=0$ case, and only for one fermion, i.e. $D=1$. (The $D=2$ case is treated in appendix C).
@@ 3 Neutrality, Dyson’s crank, and partial theta functions
To see the high-temperature behaviour of $Z_{\text{charged}}$ we need to study it as $\tau\rightarrow i0^{+}$, where $q\rightarrow 1$ and the $q$-expansion is not very helpful. As is standard, we apply modular transformations (see appendix A) to rewrite (3.2) as
@@
$Z_{\text{neutral}}(\tau)=\oint\frac{\mathrm{d}\zeta}{2\pi i\zeta}Z(u,\tau)=\frac{q^{-\frac{1}{12}}}{\eta(\tau)}\sum_{k>0,~{}k~{}\text{odd}}(-1)^{\frac{k-1}{2}}q^{k^{2}/8}=\frac{q^{\frac{1}{24}}}{\eta(\tau)}\sum_{n\geq 0}(-1)^{n}q^{\frac{n(n+1)}{2}}\;,$ (3.9)
where we set $k=2n+1$ in the last step. This series has recently been shown to be a generating function of Dyson’s non-negative crank partition function, see [34, 35], where it is given in the equivalent form
@@ 5 Modularity and high-temperature expansion
$\int_{0}^{\infty}f(x)~{}\left(\eta(ix)\right)^{3}\mathrm{d}x=\int_{0}^{\infty}\frac{F(t)}{\cosh\sqrt{\pi t}}\mathrm{d}t\;,$ (5.14)
@@
Together with the fact that $H(1,2;-1/\tau)$ is exponentially small as $\tau\rightarrow i0^{+}$ (due to the $\hat{q}^{1/8}$ factor, where $\hat{q}=e^{-2\pi i/\tau}$), we arrive at the small-$\tau$ expansion of $H(1,2;\tau$):
@@ 6 The canonical partition function
$\frac{1}{\pi^{2}}\frac{\partial^{2}\ln Z(u,\tau)}{\partial u^{2}}|_{u=0}=(2i\hat{\tau}/\pi+1/\cos(\pi u))|_{u=0}=2i\hat{\tau}/\pi+1\;.$ (6.5)
@@
$Z_{C}(\tau)=Z_{\text{charged}}(\tau)\sqrt{-i\tau}(1+\frac{\pi(-i\tau)}{4}+\cdots)\;,$ (6.7)
@@ 7 The asymptotic density of states
In this section we would like to find the asymptotic density of states of our neutral partition function, which is the main ingredient entering the Cardy formula. It is instructive to first review the computation for the full (charged) partition function, in order to highlight the differences with the neutral case. This computation is standard and reviews can be found in many sources, e.g. [5, 6, 53]. As discussed, when $\tau\rightarrow 0$, $q\rightarrow 1$ and it is convenient to re-express $Z_{\text{charged}}(\tau)=Z(u\!=\!0,\tau)$ in terms of $\hat{\tau}=-1/\tau$, such that $\hat{q}\rightarrow 0$ (see (3.4)). We define a generating function $\rho(n)$ as
@@
$\rho_{\text{neutral}}^{(0)}=\frac{\pi}{2\sqrt{2}\cdot 6^{3/4}(n+1/12)^{3/4}}\left(-i^{\frac{1}{2}}J_{3/2}(z)\right)\;.$ (7.8)
Expanding the Bessel function for large $n$ (see Appendix A), we find:
@@ 8 Discussion
Recalling our motivation from section 2, one is of course interested in how our results might apply to the higher-spin/$\mathcal{W}$ algebra correspondence. A first aspect to note is that for a comparison with the bulk one needs $D$ fermions, and although in that case the grand canonical partition function is just $Z(u,\tau)^{D}$, the restriction to the $\zeta^{0}$ part as in (3.8) is not immediate. In Appendix C we work it out for $D=2$, but ideally we would require an expression for large $D$ to make contact with the bulk. The generalisation, at $D=1$, of the non-negative crank partition function to include the spin-3 chemical potential, and which has the same relation to the partition function (2.1) as $Z_{\text{neutral}}$ does to (3.1), appears straightforward. However, one would then need to establish a suitable generalisation of the modular transformation (5.10) in order to make progress.
@@
A recent line of study [55, 56, 57] considers the modular properties of Generalised Gibbs Ensembles for real ($c=\frac{1}{2}$) free fermions with insertions of higher KdV charges, which are what would correspond to the higher $\mathcal{W}$ charges in our complex fermion case. Although the details are different, one would expect that the techniques used in those works, as well as their physical interpretation of the modular transformed equations, could provide useful input in the higher-spin context as well.
@@
Note added in v2: We have been informed [58] that, for the $D=1$ case, it is possible to go beyond our asymptotic formula (7.14) and obtain a convergent formula for the coefficients of $Z_{\text{neutral}}$. More specifically, with the numerator of $Z_{\text{neutral}}$ appropriately understood as a unary false theta function, and given that the overall weight of $Z_{\text{neutral}}$ is zero, the techniques developed in [59] allow the computation of an exact Rademacher-type formula. This can be truncated to a finite series for numerical evaluation, with the precision increasing with the number of terms included. The resulting series expressions can be seen to agree with the exact coefficients $a(n)$ to far greater accuracy than the asymptotic formula, even for very low values of $n$. As first discussed in the IIB string theory context in [60] (see [6] for a review and further references), having a full Rademacher-type expansion instead of just an asymptotic one provides a much more detailed understanding of the subleading bulk contributions to the gravitational path integral. One could therefore hope for a similar understanding to eventually emerge in our higher-spin gravity context as well. For that it would be critical to understand whether similar exact expressions are possible for the higher-$D$ cases, for which perhaps the extensions of [59] to higher rank and depth [61, 62] will be of relevance. We intend to come back to these questions in future work.
@@ Acknowledgements
We are thankful to Sam van Leuven for very enlightening discussions, and to Caner Nazaroglu for very helpful and far-reaching communications regarding our work. MM was supported by a DST/NRF Innovation Postdoctoral Fellowship during the initial stages of this work, and by the National Institute for Theoretical and Computational Sciences (NITheCS) during the later stages.
@@ Appendix A Definitions
$\theta_{1}^{\prime}(u|\tau)=\frac{\partial}{\partial u}\theta_{1}(u|\tau)=2\pi q^{\frac{1}{8}}\sum_{k=0}(-1)^{k}(2k+1)\cos(\pi(2k+1)u)q^{\frac{k(k+1)}{2}}\;.$ (A.5)
@@
$\theta_{1}^{\prime}$ is also modular, of weight $3/2$. We will only need $\theta_{1}^{\prime}(0,\tau)$, whose modular transformation follows from the equality
$\theta_{1}^{\prime}(0,\tau)=2\pi\left(\eta(\tau)\right)^{3}\;.$ (A.9)
@@ Appendix B Free fermions
For completeness, in this appendix we record the relevant generators of $\mathcal{W}_{1+\infty}$ [19], in its fermionic realisation (see e.g. [63] for a review of the various realisations of $\mathcal{W}_{1+\infty}$ and [64] for more details on the computation of the character formula). More details can be found in [17], from whose conventions we differ just by simple rescalings to simplify the presentation. We consider a complex fermion with OPE
@@ Appendix C The $D=2$ case
In section 3 we saw how the restriction of $Z(u,\tau)$ to the neutral sector works for a single fermion, $D=1$. To get a flavour of the complications that will arise in the general case, in this appendix we derive the corresponding result for $D=2$.
@@
$Z_{\text{neutral}}^{D=2}=\oint\frac{\mathrm{d}\zeta}{2\pi i\zeta}(Z(u,\tau))^{2}=1+4q^{2}+8q^{3}+13q^{4}+24q^{5}+46q^{6}+80q^{7}+139q^{9}+\cdots$ (C.1)
@@
$\begin{split}\theta_{2}^{2}(u|\tau)&=\frac{1}{2}\left(\theta_{3}(0|\tau/2)\theta_{3}(u|\tau/2)-\theta_{4}(0|\tau/2)\theta_{4}(u|\tau/2)\right)\\ &=\frac{1}{2}\left(\sum_{k\in\mathbb{Z}}q^{\frac{k^{2}}{4}}\sum_{\ell\in\mathbb{Z}}\zeta^{\ell}q^{\frac{\ell^{2}}{4}}-\sum_{k\in\mathbb{Z}}(-1)^{k}q^{\frac{k^{2}}{4}}\sum_{\ell\in\mathbb{Z}}(-1)^{\ell}\zeta^{\ell}q^{\frac{\ell^{2}}{4}}\right)\\ &=\frac{1}{2}\sum_{k\in\mathbb{Z}}\sum_{\ell\in\mathbb{Z}}\left(1-(-1)^{k+\ell}\right)\zeta^{\ell}q^{\frac{k^{2}+\ell^{2}}{4}}\\ &=\sum_{k,\ell\in\mathbb{Z},k+\ell\;\text{odd}}\zeta^{\ell}q^{\frac{k^{2}+\ell^{2}}{4}}\;.\end{split}$ (C.2)
@@
$\theta_{2}^{2}(u|\tau)=\sum_{n>0,n\;\text{odd}}(2+\zeta^{n}+\zeta^{-n})q^{\frac{n^{2}}{4}}+2\sum_{\begin{subarray}{c}m>0,n>m\\ m+n\;\text{odd}\end{subarray}}(\zeta^{m}+\zeta^{-m}+\zeta^{n}+\zeta^{-n})q^{\frac{m^{2}+n^{2}}{4}}\;.$ (C.3)
@@
$\begin{split}Z_{\text{neutral}}^{D=2}&=\frac{q^{-\frac{1}{4}}}{(q;q)_{\infty}^{2}}\oint\frac{\mathrm{d}\zeta}{2\pi i\zeta}\left(\frac{\theta_{2}(u|\tau)}{2\cos\pi u}\right)^{2}=\frac{1}{(q;q)_{\infty}^{2}}\sum_{\begin{subarray}{c}m\geq 0,n>m\\ m+n\;\text{odd}\end{subarray}}(2-\delta_{m,0})(-1)^{m}(n-m)q^{\frac{m^{2}+n^{2}-1}{4}}\\ &=\frac{1}{(q;q)_{\infty}^{2}}\left(1\!-\!2q\!+\!3q^{2}\!+\!2q^{3}\!-\!6q^{4}\!+\!3q^{6}\!+\!6q^{7}\!-\!10q^{9}\!+\!2q^{10}\!-\!6q^{11}\!+\!7q^{12}\!+\!10q^{13}+\cdots\right)\;.\end{split}$ (C.5)
@@
This slightly peculiar $q$-series agrees, after expanding the prefactor, with the empirical result (C.1). Although in this case we are not aware of any mathematical results regarding its modular properties, the physical setting is precisely the same as the $D=1$ case, so we can expect that (a) its leading behaviour for $\tau\rightarrow i0^{+}$ will be the same as the unconstrained partition function, i.e. $\sim e^{\frac{\pi i}{6\tau}-\frac{\pi i\tau}{3}}$ (just the square of the $D=1$ case), and more importantly (b) that $\hat{H}(\tau):=(q;q)_{\infty}^{2}Z_{\text{neutral}}^{D=2}$ has a well-defined expansion perturbatively in $\tau$, analogous to (5.3). If we further assume that, as for $D=1$, the restriction to the neutral sector does not change the weight [^6] This assumption is clearly not true in general, for instance restricting $\theta_{3}(u|\tau)$ to the neutral sector one would just get a constant. It is motivated by the physical expectation that the $D=2$ neutral CFT partition function would have weight zero overall, as for $D=1$. The expected small-$\tau$ expansion of $\hat{H}$ does not rely on this assumption, as it would come from the anomalous modular term., this would lead to a quantum modular form of weight 1. Of course, the first statement more or less follows from the $1/(q;q)_{\infty}^{2}$ factors (expressed in terms of $\eta(\tau)^{2}$ and modular transformed), as the remainder is unlikely to alter the exponential behaviour. Evaluating the series for large $n$, we indeed find very good numerical agreement with the leading behaviour
$\rho_{\text{neutral}}^{D=2}(n)\sim\frac{e^{\sqrt{\frac{2}{3}}\pi\sqrt{2(n+1/6)}}}{16\sqrt{3}(n+1/6)}\;.$ (C.6)
@@
$Z_{\text{neutral}}^{D=2}=\frac{e^{\frac{\pi i}{6\tau}-\frac{\pi i\tau}{3}}}{4\sqrt{2}}\sqrt{-i\tau}\left(1+\frac{\pi(-i\tau)}{4}+\cdots\right)\;,$ (C.7)